Given z1=2<45degrees, z2=3<120degrees, z3=4<180degrees. Determine the following:

a) (z1)^2+z2/z2+z3

b) z1/z2*z3


1
Expert's answer
2022-06-05T10:15:31-0400

(a)


"(z_1)^2=(2)^2\angle(2\cdot45\degree)=4i""z_2+z_3=3(-\dfrac{1}{2}+\dfrac{\sqrt{3}}{2}i)-4=-\dfrac{11}{2}-\dfrac{3\sqrt{3}}{2}i)""\dfrac{z_2}{z_2+z_3}=\dfrac{-3+3\sqrt{3}i}{-11-3\sqrt{3}i}""=\dfrac{(-3+3\sqrt{3}i)(-11-3\sqrt{3}i)}{121+27}""=\dfrac{33+9\sqrt{3}i-33\sqrt{3}i+27}{148}""=\dfrac{15}{37}-\dfrac{6\sqrt{3}}{37}i""(z_1)^2+\dfrac{z_2}{z_2+z_3}=4i+\dfrac{15}{37}-\dfrac{6\sqrt{3}}{37}i""=\dfrac{15}{37}+\dfrac{148-6\sqrt{3}}{37}i"

(b)



"z_2z_3=3(4)\angle(120\degree+180\degree)=6-6\sqrt{3}i""\dfrac{z_1}{z_2z_3}=\dfrac{\sqrt{2}+\sqrt{2}i}{6-6\sqrt{3}i}""=\dfrac{(\sqrt{2}+\sqrt{2}i)(6+6\sqrt{3}i)}{36+108}""=\dfrac{6\sqrt{2}+6\sqrt{6}i+6\sqrt{2}i-6\sqrt{6}}{144}""=-\dfrac{\sqrt{6}-\sqrt{2}}{24}+\dfrac{\sqrt{6}+\sqrt{2}}{24}i"